# Efficient and Generalizable Certified Unlearning: A Hessian-free Recollection Approach

### v3

@@ Appendix B Proofs
$\mathbf{w}_{e,b+1}^{-u}$ $\leftarrow\mathbf{w}_{e,b}^{-u}-\frac{\eta_{e,b}}{|\mathcal{B}_{e,b}|}\sum_{i\in\mathcal{B}_{e,b}}\nabla\ell\left(\mathbf{w}_{e,b}^{-u};z_{i}\right),$ (17)
@@
$\mathbf{w}_{e,b+1}$ $\leftarrow\mathbf{w}_{e,b}-\frac{\eta_{e,b}}{|\mathcal{B}_{e,b}|}\sum_{i\in\mathcal{B}_{e,b}}\nabla\ell\left(\mathbf{w}_{e,b};z_{i}\right).$
When the update includes sample $u$ during $e$-th epoch’s $b(u)$-th update, consider the linear scaling rule proposed in [28], i.e., the retraining stepsize $\hat{\eta}_{e,b(u)}=\eta_{e,b(u)}\frac{|\mathcal{B}_{e,b(u)}\backslash\{u\}|}{|\mathcal{B}_{e,b(u)}|}$, we have,
$\mathbf{w}_{e,b(u)+1}^{-u}$ $\leftarrow\mathbf{w}_{e,b(u)}^{-u}-\frac{\eta_{e,b(u)}}{|\mathcal{B}_{e,b(u)}|}\sum_{i\in\mathcal{B}_{e,b(u)}\backslash\{u\}}\nabla\ell(\mathbf{w}_{e,b(u)}^{-u};z_{i}),$ (18)
$\mathbf{w}_{e,b(u)+1}$ $\leftarrow\mathbf{w}_{e,b(u)}-\frac{\eta_{e,b(u)}}{|\mathcal{B}_{e,b(u)}|}\sum_{i\in\mathcal{B}_{e,b(u)}}\nabla\ell\left(\mathbf{w}_{e,b(u)};z_{i}\right).$
Consider the Taylor expansion of $\nabla\ell(\mathbf{w}_{e,b}^{-u};z_{i})$ around $\mathbf{w}_{e,b}$, we have,
$\nabla\ell(\mathbf{w}_{e,b}^{-u};z_{i})=\nabla\ell(\mathbf{w}_{e,b};z_{i})+\nabla^{2}\ell(\mathbf{w}_{e,b};z_{i})(\mathbf{w}_{e,b}^{-u}-\mathbf{w}_{e,b})+{o(\mathbf{w}_{e,b}^{-u}-\mathbf{w}_{e,b})},$ (19)
where $\mathbf{H}_{e,b}\!\!=\!\!\sum_{i\in\mathcal{B}_{e,b}}\!\!\!\!\nabla^{2}\ell(\mathbf{w}_{e,b};z_{i})$ denote the Hessian of the loss during the $e$-th epoch’s $b$-th update. Therefore, for (17) and (18), we can approximate the SGD recursion as,
$\mathbf{w}_{e,b+1}^{-u}-\mathbf{w}_{e,b+1}$ $\approx(\mathbf{I}-\frac{\eta_{e,b}}{|\mathcal{B}_{e,b}|}\mathbf{H}_{e,b})(\mathbf{w}_{e,b}^{-u}-\mathbf{w}_{e,b}),$ (20)
$\mathbf{w}_{e,b(u)+1}^{-u}-\mathbf{w}_{e,b(u)+1}$ $\approx\ \underbrace{(\mathbf{I}-\frac{\eta_{e,b(u)}}{|\mathcal{B}_{e,b(u)}|}\mathbf{H}_{e,b(u)})(\mathbf{w}_{e,b(u)}^{-u}-\mathbf{w}_{e,b(u)})}_{\text{Approximate impact of $u$ in previous epoches}}+\underbrace{\frac{\eta_{e,b(u)}}{|\mathcal{B}_{e,b(u)}|}\nabla\ell(\mathbf{w}_{e,b(u)};u).}_{\text{Approximate impact of $u$ in $e$-th epoch }}$
Therefore, during the $E$-th epoch, the difference between the retraining model and the learning model is approximately as,
$\mathbf{w}_{E,B}^{-u}-\mathbf{w}_{E,B}$ $\approx\prod_{b=0}^{B-1}(\mathbf{I}-\frac{\eta_{E,b}}{|\mathcal{B}_{E,b}|}\mathbf{H}_{E,b})(\mathbf{w}_{E,0}^{-u}-\mathbf{w}_{E,0})$ (21)
@@
$+\prod_{b=b(u)}^{B-1}(\mathbf{I}-\frac{\eta_{E,b}}{|\mathcal{B}_{E,b}|}\mathbf{H}_{E,b})\frac{\eta_{E,b(u)}}{|\mathcal{B}_{E,b(u)}|}\nabla\ell(\mathbf{w}_{E,b(u)};u).$
We have thus obtained the difference between the retraining model and the learning model in the $E$-th epoch. It is worth noting that the second term represents the influence of sample $u$ on the training trajectory during the $E$-th epoch, and it can be computed entirely. In the first term, $\mathbf{w}_{E,0}^{-u}-\mathbf{w}_{E,0}=\mathbf{w}_{E-1,B}^{-u}-\mathbf{w}_{E-1,B}$ is the result of the $E-1$-th epoch, and for the $E-1$ epoch, we also have a similar result. Since the initial models are identical, apply it recursively to complete the proof and, we ultimately obtain the approximator for $u$.
$\mathbf{w}_{E,B}^{-u}-\mathbf{w}_{E,B}\approx\sum_{e=0}^{E}\mathbf{M}_{e,b(u)}\nabla\ell(\mathbf{w}_{e,b(u)};u):=\mathbf{a}_{E,B}^{-u}.$ (22)
where $\mathbf{M}_{e,b(u)}=\frac{\eta_{e,b(u)}}{|\mathcal{B}_{e,b(u)}|}\prod^{E}_{k=e}\prod_{b=b(u)+1}^{B-1}(\mathbf{I}-\frac{\eta_{k,b}}{|\mathcal{B}_{k,b}|}\mathbf{H}_{k,b})$. ∎
B.2 Detailed Proof of Theorem 1
@@
Without loss of generality, we assume that $b(u_{1})<b(u_{2})$. Based on (10), the approximators for $u_{1}$ and $u_{2}$ are denoted as $\mathbf{a}_{E,B}^{-u_{1}}$ and $\mathbf{a}_{E,B}^{-u_{2}}$, respectively, as follows:
$\mathbf{a}_{E,B}^{-u_{1}}$ $=\sum_{e=0}^{E}\mathbf{M}_{e,b(u_{1})}\nabla\ell(\mathbf{w}_{e,b(u_{1})};u_{1}),$ (23)
$\mathbf{a}_{E,B}^{-u_{2}}$ $=\sum_{e=0}^{E}\mathbf{M}_{e,b(u_{2})}\nabla\ell(\mathbf{w}_{e,b(u_{2})};u_{2}).$
When we simultaneously delete $u_{1}$ and $u_{2}$, the difference between retraining the model and learning the model is as follows,
@@
$\mathbf{w}_{E,B}^{-u_{2}-u_{1}}-\mathbf{w}_{E,B}$ $\approx\prod_{b=0}^{B-1}(\mathbf{I}-\frac{\eta_{E,b}}{|\mathcal{B}_{E,b}|}\mathbf{H}_{E,b})(\mathbf{w}_{E,0}^{-u_{2}-u_{1}}-\mathbf{w}_{E,0})$ (27)
$+\prod_{b=b(u_{1})}^{B-1}(\mathbf{I}-\frac{\eta_{E,b}}{|\mathcal{B}_{E,b}|}\mathbf{H}_{E,b})\frac{\eta_{E,b(u_{1})}}{|\mathcal{B}_{E,b(u_{1})}|}\nabla\ell(\mathbf{w}_{E,b(u_{1})};u_{1})$
$+\prod_{b=b(u_{2})}^{B-1}(\mathbf{I}-\frac{\eta_{E,b}}{|\mathcal{B}_{E,b}|}\mathbf{H}_{E,b})\frac{\eta_{E,b(u_{2})}}{|\mathcal{B}_{E,b(u_{2})}|}\nabla\ell(\mathbf{w}_{E,b(u_{2})};u_{2}).$
Apply it recursively to complete the proof, and we can simultaneously compute the impact of $u_{1}$ and $u_{2}$ on training for all epochs. Since the initial models are identical, recursion stops at $e=0,b=b(u_{1})$, and we eventually obtain,
$\mathbf{w}_{E,B}^{-u_{2}-u_{1}}-\mathbf{w}_{E,B}\approx\sum_{e=0}^{E}\mathbf{M}_{e,b(u_{1})}\nabla\ell(\mathbf{w}_{e,b(u_{1})};u_{1})+\sum_{e=0}^{E}\mathbf{M}_{e,b(u_{2})}\nabla\ell(\mathbf{w}_{e,b(u_{2})};u_{2}):=\mathbf{a}_{E,B}^{-u_{2}-u_{1}}.$ (28)
According to (23) and (28), we have completed the proof of $\mathbf{a}_{E,B}^{-u_{2}-u_{1}}=\mathbf{a}_{E,B}^{-u_{1}}+\mathbf{a}_{E,B}^{-u_{2}}$.
B.3 Detailed Proof of Lemma 2
@@
Now we begin the proof that the remainder term $o(\mathbf{w}_{e,b}^{-u}-\mathbf{w}_{e,b})$ in the $e$-th epoch, $b$-th round is bounded by gradient clipping. Specifically, recalling (4), we have,
@@
In model training, we threshold the stochastic gradient norm $\|\nabla\ell(\mathbf{w}_{e,b};u)\|$ at $C$ when clipping is applied.
$\|\mathbf{w}_{e,b}^{-u}-\mathbf{w}_{e,b}\|$ $\leq\|\mathbf{w}_{e,b-1}^{-u}-\mathbf{w}_{e,b-1}\|+2\eta_{e,b-1}C.$ (30)
Consider a simple stepsize time decay strategy $\eta_{e,b}=q\eta_{e,b-1}$, where $10<q<$ 1 is the decay rate, and let $\eta$ be the initial stepsize. Recalling that $t=eB+b$, then we get
$\|\mathbf{w}_{e,b}^{-u}-\mathbf{w}_{e,b}\|\leq\|\mathbf{w}_{e,b-1}^{-u}-\mathbf{w}_{e,b-1}\|+2\eta Cq^{t-1}.$ (31)
@@
$\|\mathbf{w}_{e,b}^{-u}-\mathbf{w}_{e,b}\|$ $\leq 2\eta Cq^{t-1}+2\eta Cq^{t-2}+...+2\eta Cq^{0}$ (32)
$=2\eta C\sum_{k=0}^{t-1}q^{k}$
$=2\eta C\frac{q^{t}-1}{q-1}.$
Therefore, we have completed the proof. ∎
B.4 Detailed Proof of Theorem 4
@@
Let us recall the approximator of (10) in Section 3, we have,
$\mathbf{a}_{E,B}^{-u}=$ $\sum_{e=1}^{E}\mathbf{M}_{e,b(u)}\nabla\ell(\mathbf{w}_{e,b(u)};u),$ (33)
$\text{where}\ \mathbf{M}_{e,b(u)}=$ $\frac{\eta_{e,b(u)}}{|\mathcal{B}_{e,b(u)}|}\prod^{E}_{k=e}\prod_{b=b(u)+1}^{B-1}\Big{(}\mathbf{I}-\frac{\eta_{k,b}}{|\mathcal{B}_{k,b}|}\mathbf{H}_{k,b}\Big{)}.$
@@
$\mathbf{w}_{E,B}^{-u}-\mathbf{w}_{E,B}-\mathbf{a}_{E,B}^{-u}=$ $(\mathbf{I}-\frac{\eta_{E,B-1}}{|\mathcal{B}_{E,B-1}|}\mathbf{H}_{E,B-1})(\mathbf{w}_{E,B-1}^{-u}-\mathbf{w}_{E,B-1})$ (34)
@@
For convenience, we define ${\Delta^{-u}_{E,B}}=\mathbf{w}_{E,B}^{-u}-\mathbf{w}_{E,B}$, and $\hat{\mathbf{H}}_{E,B-1}=\mathbf{I}-\frac{\eta_{E,B-1}}{|\mathcal{B}_{E,B-1}|}\mathbf{H}_{E,B-1}$. Here, we begin to analyze the error generated by each update,
$\Delta^{-u}_{E,B}-\mathbf{a}_{E,B}^{-u}$ $=\hat{\mathbf{H}}_{E,B-1}\Delta^{-u}_{E,B-1}+\eta_{E,B-1}o(\Delta^{-u}_{E,B-1})-\mathbf{a}_{E,B}^{-u},$ (35)
Furthermore, we observe that during the $E$-th epoch, (35) can be obtained from affine stochastic recursion as follows,
$\|\Delta^{-u}_{E,B}-\mathbf{a}_{E,B}^{-u}\|$ $\leq\|\prod_{b=0}^{B-1}\hat{\mathbf{H}}_{E,b}\Delta^{-u}_{E,0}-\mathbf{a}_{E-1,B}^{-u}\|$ (36)
@@
$+\underbrace{\eta q^{T-1}\|\Delta^{-u}_{E,B-1}\|+\eta q^{T-2}\|\hat{\mathbf{H}}_{E,B-1}\Delta^{-u}_{E,B-2}\|+...+\eta q^{T-B}\|\prod_{b=0}^{B-1}\hat{\mathbf{H}}_{E,b}\Delta^{-u}_{E,0}\|.}_{\text{Approximation error term$C_{1}$ during the $E$-th epoch}}$
@@
$\|\Delta^{-u}_{e,b}\|=\|\mathbf{w}_{e,b}^{-U}-\mathbf{w}_{e,b}\|\leq 2\eta C\frac{1-q^{t}}{1-q}.$ (37)
Recalling Lemma 3, we have that for all $e,b$ and $\rho$ is the spectral radius of $\mathbf{I}-\frac{\eta_{e,b}}{|\mathcal{B}_{e,b}|}\mathbf{H}_{e,b}$,
$\|\hat{\mathbf{H}}_{e,b}\Delta^{-u}_{e,b}\|=\|(\mathbf{I}-\frac{\eta_{e,b}}{|\mathcal{B}_{e,b}|}\mathbf{H}_{e,b})\Delta^{-u}_{e,b}\|\leq\rho\|\Delta^{-u}_{e,b}\|\leq\rho\cdot 2\eta C\frac{1-q^{t}}{1-q},$ (38)
@@
$q^{T-1}$ $\|\Delta^{-u}_{E,B-1}\|+q^{T-2}\|\hat{\mathbf{H}}_{E,B-1}\Delta^{-u}_{E,B-2}\|+...+q^{T-B}\|\prod_{b=0}^{B-1}\hat{\mathbf{H}}_{E,b}\Delta^{-u}_{E,0}\|$ (39)
$\leq 2\eta C\frac{q^{T-1}-q^{2(T-1)}}{1-q}+2\eta C\frac{q^{T-2}-q^{2(T-2)}}{1-q}\rho+...+2\eta C\frac{q^{T-B}-q^{2(T-B)}}{1-q}\rho^{B-1}$
$=\frac{2\eta C}{1-q}\sum_{k=T-1}^{T-B}\rho^{T-1-k}(q^{k}-q^{2k}).$
(2) we then bound the second approximation error term $C_{2}$ during the $E$-th epoch. We threshold the stochastic gradient norm $\|\nabla\ell(\mathbf{w}_{e,b};u)\|$ at $C$ when clipping is applied, and we thus have the following upper bound,
$\frac{\eta q^{T-B+b(u)}}{|\mathcal{B}_{E,b(u)}|}\|\prod_{b=b(u)+1}^{B-1}\hat{\mathbf{H}}_{E,b}(\nabla\ell(\mathbf{w}^{-u}_{E,b(u)};u)-\nabla\ell(\mathbf{w}_{E,b(u)};u))\|\leq\frac{2C\eta}{|\mathcal{B}|}q^{T-B+b(u)}\rho^{B-b(u)-1}.$ (40)

### v4

@@ Appendix C Proofs
$\mathbf{w}_{E,B+1}\leftarrow\mathbf{w}_{E,B}-\frac{\eta_{E,B}}{|\mathcal{B}_{E,B}|}\sum_{i\in\mathcal{B}_{E,B}}\left(\nabla\ell\left(\mathbf{w}_{E,B};z_{i}\right)+\bm{1}_{\mathrm{b=b(u)}}\omega\nabla\ell\left(\mathbf{w}_{E,B};u\right)\right).$ (26)
where $\bm{1}_{\mathrm{b=b(u)}}$ is an indicator function that equals 1 if the condition $b=b(u)$ is true (that is, sample $u$ is to be involved (sampled) in the $b(u)$-th update of the $e$-th epoch), and 0 otherwise. We define that when $\omega=0$, Eq. 26 represents the original learning update, whereas $\omega=-1$ represents the update after forgetting training sample $u$. To obtain the difference between the learning model and retraining model, we calculate the derivative of $\mathbf{w}_{E,B+1}$ w.r.t the weighting factor at $\omega=0$, given by,
$\mathbf{w}_{E,B+1}^{-u}-\mathbf{w}_{E,B+1}=\left.\mathbf{w}_{E,B+1}\right|_{\omega=-1}-\left.\mathbf{w}_{E,B+1}\right|_{\omega=0}\approx-\left.\frac{d\mathbf{w}_{E,B+1}}{d\omega}\right|_{\omega=0}.$ (27)
Based on the chain rule, we have,
$\left.\frac{d{\mathbf{w}}_{E,B+1}}{d\omega}\right|_{\omega=0}=\left.\frac{d\mathbf{w}_{E,B+1}}{d\mathbf{w}_{E,B}}\frac{d\mathbf{w}_{E,B}}{d\omega}\right|_{\omega=0}.$ (28)
Let $\mathbf{H}_{e,b}={1}/{|\mathcal{B}_{e,b}|}\sum_{i\in\mathcal{B}_{e,b}}\nabla^{2}\ell(\mathbf{w}_{e,b};z_{i})$ denote the Hessian matrix of the loss function. For Eq. 26, we have that $\frac{d\mathbf{w}_{E,B+1}}{d\mathbf{w}_{E,B}}=\mathbf{I}-{\eta_{E,B}}\mathbf{H}_{E,B}$. Substituting into Eq. 28,
$\left.\frac{d{\mathbf{w}}_{E,B+1}}{d\omega}\right|_{\omega=0}=\left.\left(\mathbf{I}-{\eta_{E,B}}\mathbf{H}_{E,B}\right)\frac{d\mathbf{w}_{E,B}}{d\omega}\right|_{\omega=0},$ (29)
which is equivalent to Eq. 8 in the main text (Case 1), that is,
$\mathbf{w}_{e,b+1}^{-u}-\mathbf{w}_{e,b+1}\approx(\mathbf{I}-{\eta_{e,b}}\mathbf{H}_{e,b})(\mathbf{w}_{e,b}^{-u}-\mathbf{w}_{e,b}).$ (30)
Recursively applying the chain rule until $\bm{1}_{\mathrm{b=b(u)}}=1$ during $E$-th epoch,
$\left.\frac{d{\mathbf{w}}_{E,B+1}}{d\omega}\right|_{\omega=0}=\prod_{b=B}^{b(u)+1}\left.\left(\mathbf{I}-{\eta_{E,b}}\mathbf{H}_{E,b}\right)\frac{d\mathbf{w}_{E,b(u)+1}}{d\omega}\right|_{\omega=0}.$ (31)
When the indicator function $\bm{1}_{\mathrm{b=b(u)}}=1$, for Eq. 26, we obtain,
$\left.\frac{d\mathbf{w}_{E,b(u)+1}}{d\omega}\right|_{\omega=0}=\left.\frac{d\mathbf{w}_{E,b(u)}}{d\omega}\right|_{\omega=0}-\eta_{E,b(u)}\mathbf{H}_{E,b(u)}\left.\frac{d\mathbf{w}_{E,b(u)}}{d\omega}\right|_{\omega=0}-\frac{\eta_{E,b}}{|\mathcal{B}_{E,b}|}\nabla\ell\left(\mathbf{w}_{E,b(u)};u\right),$ (32)
which is equivalent to Eq. 9 in the main text (Case 2), that is,
$\!\mathbf{w}_{e,b(u)+1}^{-u}\!\!-\!\mathbf{w}_{e,b(u)+1}\!\approx\!(\mathbf{I}\!-\!{\eta_{e,b(u)}}\mathbf{H}_{e,b(u)})(\mathbf{w}_{e,b(u)}^{-u}\!-\!\mathbf{w}_{e,b(u)})\!+\!\!\frac{\eta_{e,b(u)}}{|\mathcal{B}_{e,b(u)}|}\nabla\ell(\mathbf{w}_{e,b(u)};u).$ (33)
Substituting Eq. 32 into Eq. 31,
@@
$\left.\frac{d{\mathbf{w}}_{E,B+1}}{d\omega}\right|_{\omega=0}\!\!\!=\prod_{b=B}^{b(u)}\left.\left(\mathbf{I}-{\eta_{E,b}}\mathbf{H}_{E,b}\right)\frac{d\mathbf{w}_{E,b(u)}}{d\omega}\right|_{\omega=0}\!\!\!-\frac{\eta_{e,b}}{|\mathcal{B}_{e,b}|}\prod_{b=B}^{b(u)+1}\left(\mathbf{I}-{\eta_{E,b}}\mathbf{H}_{E,b}\right)\nabla\ell\left(\mathbf{w}_{E,b(u)};u\right).$ (34)
Therefore, the impact of sample $u$ on the model update trajectory during the $E$-th round is,
$\mathbf{w}_{E,B}^{-u}-\mathbf{w}_{E,B}$ $\approx\prod_{b=B-1}^{0}(\mathbf{I}-{\eta_{E,b}}\mathbf{H}_{E,b})(\mathbf{w}_{E,0}^{-u}-\mathbf{w}_{E,0})$ (35)
$+\prod_{b=B-1}^{b(u)+1}(\mathbf{I}-{\eta_{E,b}}\mathbf{H}_{E,b})\frac{\eta_{E,b(u)}}{|\mathcal{B}_{E,b(u)}|}\nabla\ell(\mathbf{w}_{E,b(u)};u).$
Since initially the retraining model $\mathbf{w}_{0,0}^{-u}$ and the learning model $\mathbf{w}_{0,0}$ are the same, the first term of Eq. 35 will recursively reduce to the initial model and will equal 0, and as sample $u$ participates in updates once per epoch, the second term is the sum over $E$ epochs. [^5] Note that we assume sample $u$ participates in the gradient calculation exactly once per epoch. Therefore, in Eq. 36, the number of summations $E$ represents that sample $u$ has participated in $E$ gradient calculations. In general, due to the randomness of sampling, sample $u$ may be sampled multiple times or not at all within a single epoch. Nevertheless, we can derive the results for these situations in the same manner. To facilitate understanding and draw relevant conclusions, all results in this work are based on the simplified assumption. Apply the above process recursively to complete the proof and, we obtain the approximator for sample $u$.
@@
${\mathbf{a}_{E,B}^{-u}:=\sum_{e=0}^{E}\mathbf{M}_{e,b(u)}\cdot\nabla\ell(\mathbf{w}_{e,b(u)};u),\ \text{where}\ \mathbf{M}_{e,b(u)}=\frac{\eta_{e,b(u)}}{|\mathcal{B}_{e,b(u)}|}\hat{\mathbf{H}}_{E,B-1\rightarrow e,b(u)+1}}.$ (36)
We define $\hat{\mathbf{H}}_{E,B-1\rightarrow e,b(u)+1}\!=\!(\mathbf{I}\!-\!{\eta_{E,B-1}}\mathbf{H}_{E,B-1})...(\mathbf{I}\!-\!{\eta_{e,b(u)+1}}\mathbf{H}_{e,b(u)+1})$, which means that the multiplication should begin with recent update and continue backward through previous updates.
C.2 Detailed Proof of Theorem 1
@@
Without loss of generality, we assume that $b(u_{1})<b(u_{2})$. Based on Eq. 11, the approximators for $u_{1}$ and $u_{2}$ are denoted as $\mathbf{a}_{E,B}^{-u_{1}}$ and $\mathbf{a}_{E,B}^{-u_{2}}$, respectively, as follows:
$\mathbf{a}_{E,B}^{-u_{1}}$ $:=\sum_{e=0}^{E}\mathbf{M}_{e,b(u_{1})}\cdot\nabla\ell(\mathbf{w}_{e,b(u_{1})};u_{1}),\ \text{where}\ \mathbf{M}_{e,b(u_{1})}=\frac{\eta_{e,b(u_{1})}}{|\mathcal{B}_{e,b(u_{1})}|}\hat{\mathbf{H}}_{E,B-1\rightarrow e,b(u_{1})+1},$ (37)
$\mathbf{a}_{E,B}^{-u_{2}}$ $:=\sum_{e=0}^{E}\mathbf{M}_{e,b(u_{2})}\cdot\nabla\ell(\mathbf{w}_{e,b(u_{2})};u_{2}),\ \text{where}\ \mathbf{M}_{e,b(u_{2})}=\frac{\eta_{e,b(u_{2})}}{|\mathcal{B}_{e,b(u_{2})}|}\hat{\mathbf{H}}_{E,B-1\rightarrow e,b(u_{2})+1}$
When we simultaneously delete $u_{1}$ and $u_{2}$, the difference is,
@@
$\mathbf{w}_{E,B}^{-u_{2}-u_{1}}-\mathbf{w}_{E,B}\approx\prod_{b=B-1}^{0}(\mathbf{I}-\frac{\eta_{E,b}}{|\mathcal{B}_{E,b}|}\mathbf{H}_{E,b})(\mathbf{w}_{E,0}^{-u_{2}-u_{1}}-\mathbf{w}_{E,0})+$ (41)
$\!\!\!\prod_{b=B-1}^{b(u_{1})+1}\!\!\!\!(\mathbf{I}\!-\!\frac{\eta_{E,b}}{|\mathcal{B}_{E,b}|}\mathbf{H}_{E,b})\frac{\eta_{E,b(u_{1})}}{|\mathcal{B}_{E,b(u_{1})}|}\!\nabla\ell(\mathbf{w}_{E,b(u_{1})};\!u_{1})\!\!+\!\!\!\!\!\!\prod_{b=B-1}^{b(u_{2})+1}\!\!\!\!(\mathbf{I}\!-\!\frac{\eta_{E,b}}{|\mathcal{B}_{E,b}|}\mathbf{H}_{E,b})\frac{\eta_{E,b(u_{2})}}{|\mathcal{B}_{E,b(u_{2})}|}\!\nabla\ell(\mathbf{w}_{E,b(u_{2})};\!u_{2}).$
Apply it recursively to complete the proof, and we can simultaneously compute the impact of $u_{1}$ and $u_{2}$ on training for all epochs.
$\mathbf{w}_{E,B}^{-u_{2}-u_{1}}-\mathbf{w}_{E,B}\approx\mathbf{a}_{E,B}^{-u}:=\sum_{e=0}^{E}\mathbf{M}_{e,b(u_{1})}\cdot\nabla\ell(\mathbf{w}_{e,b(u_{1})};u_{1})+\sum_{e=0}^{E}\mathbf{M}_{e,b(u_{2})}\cdot\nabla\ell(\mathbf{w}_{e,b(u_{2})};u_{2})+.$ (42)
According to Eqs. 37 and 42, we have completed the proof of $\mathbf{a}_{E,B}^{-u_{2}-u_{1}}=\mathbf{a}_{E,B}^{-u_{1}}+\mathbf{a}_{E,B}^{-u_{2}}$.
C.3 Detailed Proof of Lemma 2
@@
Now we begin the proof that the remainder term $o(\mathbf{w}_{e,b}^{-u}-\mathbf{w}_{e,b})$ in the $e$-th epoch, $b$-th round is bounded. Specifically, recalling Eq. 4, we have,
@@
For all $e,b$ and $z\in\mathcal{B}_{e,b}$, suppose that exists $G$ such that $G\!=\!\max\|\nabla\ell(\mathbf{w}_{e,b};z)\|\!<\!\infty$.
$\|\mathbf{w}_{e,b}^{-u}-\mathbf{w}_{e,b}\|$ $\leq\|\mathbf{w}_{e,b-1}^{-u}-\mathbf{w}_{e,b-1}\|+2\eta_{e,b-1}G.$ (44)
Consider a simple geometrically step size decay strategy $\eta_{e,b}=q\eta_{e,b-1}$, where $10<q<$ 1 is the decay rate, and let $\eta$ be the initial stepsize.666 Li et al. (2021) demonstrated that SGD with exponential stepsizes achieves (almost) optimal convergence rate for smooth non-convex functions, thus not affecting the convergence of learning process in our paper. Recalling that $t=eB+b$, then we get
$\|\mathbf{w}_{e,b}^{-u}-\mathbf{w}_{e,b}\|\leq\|\mathbf{w}_{e,b-1}^{-u}-\mathbf{w}_{e,b-1}\|+2\eta Gq^{t-1}.$ (45)
@@
$\|\mathbf{w}_{e,b}^{-u}-\mathbf{w}_{e,b}\|$ $\leq 2\eta Gq^{t-1}+2\eta Gq^{t-2}+...+2\eta Gq^{0}$ (46)
$=2\eta G\sum_{k=0}^{t-1}q^{k}$
$=2\eta G\frac{q^{t}-1}{q-1}.$
Therefore, we have completed the proof of Lemma 2. ∎
C.4 Detailed Proof of Theorem 4
@@
Let us recall the approximator of Eq. 11 in Section 3, we have,
$\mathbf{a}_{E,B}^{-u}:=\sum_{e=0}^{E}\mathbf{M}_{e,b(u)}\cdot\nabla\ell(\mathbf{w}_{e,b(u)};u),\ \text{where}\ \mathbf{M}_{e,b(u)}=\frac{\eta_{e,b(u)}}{|\mathcal{B}_{e,b(u)}|}\hat{\mathbf{H}}_{E,B-1\rightarrow e,b(u)+1}.$ (47)
@@
$\mathbf{w}_{E,B}^{-u}-\mathbf{w}_{E,B}-\mathbf{a}_{E,B}^{-u}=$ $(\mathbf{I}-{\eta_{E,B-1}}\mathbf{H}_{E,B-1})(\mathbf{w}_{E,B-1}^{-u}-\mathbf{w}_{E,B-1})$ (48)
@@
For convenience, we define ${\Delta\mathbf{w}^{-u}_{E,B}}=\mathbf{w}_{E,B}^{-u}-\mathbf{w}_{E,B}$, and $\hat{\mathbf{H}}_{E,B-1}=\mathbf{I}-{\eta_{E,B-1}}\mathbf{H}_{E,B-1}$. Here, we begin to analyze the error generated by each update,
$\Delta\mathbf{w}^{-u}_{E,B}-\mathbf{a}_{E,B}^{-u}$ $=\hat{\mathbf{H}}_{E,B-1}\Delta\mathbf{w}^{-u}_{E,B-1}+\eta_{E,B-1}o(\Delta\mathbf{w}^{-u}_{E,B-1})-\mathbf{a}_{E,B}^{-u},$ (49)
Furthermore, we observe that during the $E$-th epoch, Eq. 49 can be obtained from affine stochastic recursion as follows,
$\|\Delta\mathbf{w}^{-u}_{E,B}-\mathbf{a}_{E,B}^{-u}\|$ $\leq\|\prod_{b=0}^{B-1}\hat{\mathbf{H}}_{E,b}\Delta\mathbf{w}^{-u}_{E,0}-\mathbf{a}_{E-1,B}^{-u}\|$ (50)
@@
$+$ $\underbrace{\eta q^{T-1}\|\Delta\mathbf{w}^{-u}_{E,B-1}\|+\eta q^{T-2}\|\hat{\mathbf{H}}_{E,B-1}\Delta\mathbf{w}^{-u}_{E,B-2}\|+...+\eta q^{T-B}\|\prod_{b=0}^{B-1}\hat{\mathbf{H}}_{E,b}\Delta\mathbf{w}^{-u}_{E,0}\|.}_{\text{Approximation error term $C_{1}$ during the $E$-th epoch}}$
@@
$\|\Delta\mathbf{w}^{-u}_{e,b}\|=\|\mathbf{w}_{e,b}^{-U}-\mathbf{w}_{e,b}\|\leq 2\eta G\frac{1-q^{t}}{1-q}.$ (51)
Recalling Lemma 3, we have that for all $e,b$ and $\rho$ is the spectral radius of $\mathbf{I}-{\eta_{e,b}}\mathbf{H}_{e,b}$,
$\|\hat{\mathbf{H}}_{e,b}\Delta\mathbf{w}^{-u}_{e,b}\|=\|(\mathbf{I}-{\eta_{e,b}}\mathbf{H}_{e,b})\Delta\mathbf{w}^{-u}_{e,b}\|\leq\rho\|\Delta\mathbf{w}^{-u}_{e,b}\|\leq\rho\cdot 2\eta G\frac{1-q^{t}}{1-q},$ (52)
@@
$q^{T-1}$ $\|\Delta\mathbf{w}^{-u}_{E,B-1}\|+q^{T-2}\|\hat{\mathbf{H}}_{E,B-1}\Delta\mathbf{w}^{-u}_{E,B-2}\|+...+q^{T-B}\|\prod_{b=0}^{B-1}\hat{\mathbf{H}}_{E,b}\Delta\mathbf{w}^{-u}_{E,0}\|$ (53)
$\leq 2\eta G\frac{q^{T-1}-q^{2(T-1)}}{1-q}+2\eta G\frac{q^{T-2}-q^{2(T-2)}}{1-q}\rho+...+2\eta G\frac{q^{T-B}-q^{2(T-B)}}{1-q}\rho^{B-1}$
$=\frac{2\eta G}{1-q}\sum_{k=T-1}^{T-B}\rho^{T-1-k}(q^{k}-q^{2k}).$
(2) we then bound the second approximation error term $C_{2}$ during the $E$-th epoch. Suppose the stochastic gradient norm $\|\nabla\ell(\mathbf{w}_{e,b};u)\|$ is bounded by $G$, and we thus have the following upper bound,
$\frac{\eta q^{T-B+b(u)}}{|\mathcal{B}_{E,b(u)}|}\|\prod_{b=b(u)+1}^{B-1}\hat{\mathbf{H}}_{E,b}(\nabla\ell(\mathbf{w}^{-u}_{E,b(u)};u)-\nabla\ell(\mathbf{w}_{E,b(u)};u))\|\leq\frac{2G\eta}{|\mathcal{B}|}q^{T-B+b(u)}\rho^{B-b(u)-1}.$ (54)